# Pebble Games and Algebraic Proof Systems

### v1

@@ 3 Monomial Calculus and pebbling formulas
Proof.
It is convenient to consider here the equivalent notion of white pebbling. Let $\mathcal{P}=(\mathbb{P}_{0},\dots,\mathbb{P}_{t})$ be a white pebbling strategy for $G$ with $\mathbb{P}_{1}=\{z\}$ and $\mathbb{P}_{t}=\emptyset$ using $s$ pebbles. We show that for each pebbling configuration $\mathbb{P}_{i}$, $i\in[t]$, $\mathbb{P}_{i}=\{v_{i_{1}},\dots,v_{i_{k_{i}}}\}$ the monomial $m_{i}=\prod_{v\in\mathbb{P}_{i}}x_{v}$ can be derived from $\mathrm{Peb}_{G}$ and $m_{i-1}$ in degree $s$ and size 1 if $\mathbb{P}_{i}$ adds a pebble, or size $12s-$ 1 if $\mathbb{P}_{i}$ removes a pebble. This proves the result since in the $t$ steps of the pebbling strategy half of the steps add a pebble and the other half of the steps remove a pebble (each added pebble has to be removed). The total number of steps is therefore $\frac{t}{2}+\frac{t}{2}(2s-1)=ts.$
@@
Pebble removal: If the configuration at pebbling step $i+1$ is reached after removing a white pebble from vertex $v$ and $\mathbb{P}_{i}=\{v,u_{i_{1}},\dots,u_{i_{k_{i}}}\}$ with $k_{i}\leq s-1$ then all predecessors of $v$ are in the set $\{u_{i_{1}},\dots,u_{i_{k_{i}}}\}.$ For the derivation of $m_{i+1}$ we can multiply the axiom $(1-x_{v})\prod_{u\in\operatorname{\mathrm{p}red}(v)}x_{u}$ by the variables in $\mathrm{Var}(m_{i})\setminus(\bigcup_{u\in\operatorname{\mathrm{p}red}(v)}x_{u}\cup\{x_{v}\})$, and add this polynomial to $m_{i}$ obtaining $m_{i+1}$. Since there are at most $s-1$ variables in $\mathrm{Var}(m_{i})\setminus(\bigcup_{u\in\operatorname{\mathrm{p}red}(v)}x_{u}\cup\{x_{v}\})$, the number of intermediate monomials added to the proof (counting also monomial $m_{i+1}$) is at most $12(s-1)+1=2s-$ 1 .
Observe that in all the steps in the refutation, at most two different monomials are active and the number of different variables in these monomials coincides with the largest of their degrees. This shows that the variable space of the MC refutation is also bounded by $s$. ∎
Observation 10.
The size bound $ts$ in the above proof comes from the way the MC rules are defined. As is the case of PC, in the multiplication rule only one variable at at time is multiplied, even when multiplying the axiom polynomials. When an axiom is multiplied by a monomial with several variables, all the intermediate polynomials contribute to the size of the MC refutation. This is different from the the usual way to measure the size in the NS case, where intermediate monomials are not counted. If we would define the MC rules as those in NS, that is, if a whole monomial could be multiplied by an axiom in one step, the size of the MC proof would be would avoid the $s$ factor in the monomial size and obtain size $2t$ instead.
In order to prove a result in the other direction we consider a very restricted kind of refutation in MC, analogous to what is known as an input refutation in Resolution. It this kind of refutation in every Resolution step one of the parent clauses must be an axiom. Input Resolution is not complete, but it is complete for Horn formulas. We will show that the same is true for MC input refutations.
Definition 11.
A MC refutation $\pi$ of a contradictory set of polynomials $F$ is called an input refutation if there is a sequence of monomials $M_{0},\dots,M_{t}$ such that $M_{0}$ is the product of a monomial and an axiom, $M_{t}=1$ and for each $i$ $M_{i}$ is obtained by multiplying $M_{i-1}$ times a variable, or by the linear combination rule from $M_{i-1}$ and a monomial multiplied by an axiom polynomial. We will call the sequence of monomials $M_{0},\dots,M_{t}$ the backbone of the proof.
@@
Proof.
Let $d$ and $t$ be the degree and size of $\pi$. We can suppose that $\pi$ is multilinear. We prove the result by induction on $k$, the number of times the multiplication rule is applied to a monomial derived in $\pi$. In the base case $k=0$, $\pi$ is just a NS refutation of $P_{F}.$ This means that there is a linear combination of a set of polynomials $S$ that adds up to 1. Each of these polynomials has the form of a polynomial axiom multiplied by a monomial and since $F$ is a Horn formula, each polynomial in $S$ has either one or two monomials with coefficients in $\{1,-1\}$. We will represent such a polynomial $p=\alpha_{m}m+\alpha_{m^{\prime}}m^{\prime}$ by the pair of monomials $(m,m^{\prime})$. Since we are representing monomials by their set of variables, the monomial $1$ is represented by $\emptyset$. Clauses without positive literals are encoded as single monomials $(m)$. Some polynomial in $S$ has a single monomial otherwise the whole set $S$ would have a common root by setting all variables to 1. Moreover, we claim there has to be a sequence of polynomials $p_{0},\dots,p_{\ell}$ in $S$ represented by the monomials $(\emptyset,m_{1}),(m_{1},m_{2}),(m_{2},m_{3})\dots,(m_{\ell-1},m),(m).$ In order to see this, we define an undirected graph whose vertices are the different monomials in $S$ and there is an edge between two monomials $m,m^{\prime}$ if both appear in the same polynomial in $S$. In a minimal $\mathrm{NS}$ proof we only need to consider the connected component containing monomial $1$. Such component must have a vertex $v$ coming from a polynomial with a single monomial $m$ in $S$ because any set of binomials in $S$ can be satisfied by assigning all variables to 1. Therefore there is a path in the graph connecting vertices $1$ and $v$. Such a path with minimal length defines a sequence of edges $(1,v_{1}),(v_{1},v_{2}),\dots,(v_{\ell-1},v)$, and from it we can build the chain of polynomials $(\emptyset,m_{1}),(m_{1},m_{2}),(m_{2},m_{3})\dots,(m_{\ell-1},m),(m)$. This proves the claim.
@@
For the case $k>0$ let $m^{\prime}$ be the first monomial in the proof that is the result of a multiplication from a derived monomial $m$ and a variable $x$, $m^{\prime}=xm$ in $\pi$. The same argument as above shows that there is a sequence of polynomials $p_{0},\dots,p_{\ell},\hat{m}$ in $\pi$ from which we can extract an input monomial refutation that starts at $M_{0}=\hat{m}$ and derives at some point $M_{i}=m$. In the next step the multiplication rule is applied to obtain $M_{i+1}=m^{\prime}.$ Observe that the set of polynomials $m^{\prime}\cup P_{F}$ still has the Horn property and that there is a sub-proof of $\pi$ that derives the monomial 1 from this set applying the multiplication rule at most $k-1$ times. By induction hypothesis we know that there is a sequence of polynomials $p^{\prime}_{0},\dots,p^{\prime}_{\ell^{\prime}}$ in $\pi$ represented by the monomials $(\emptyset,m^{\prime}_{1}),(m^{\prime}_{1},m^{\prime}_{2}),\dots,(m^{\prime}_{r},m^{\prime})$ from which an input refutation of $P_{F}\cup m^{\prime}$ can be extracted. We can put together both input MC refutations $M_{0}\dots M_{i}$ and $M_{i+1},\dots,1$. Again we can assume that all the monomials in the backbone are different since otherwise we could shorten $\pi^{\prime}$. By the same argument as in the base case the size and degree of the input MC refutation cannot be larger than that of $\pi.$
@@
Because of Corollary 13 we can suppose that there exists an input MC refutation with monomials $M_{0},\dots M_{t}$ starting with $M_{1}=mx_{\rm sink}$ for some monomial $m$ and with $M_{t}=1$. We describe a strategy for a white pebbling of $G$ following $\pi$. As explained in the preliminaries, this is equivalent to a black pebbling strategy. At each step $i$ only the vertices corresponding to variables in $M_{i}$ have a pebble on them. In a multiplication step a new pebble is added, which is always possible in a white pebbling strategy. We only have to show that in case variables disappear when going from $M_{i}$ to $M_{i+1}$, this is also a correct pebbling move. But in this case, the step from $i$ to $i+1$ is a linear combination of $M_{i}$ with the axiom for some vertex $v$, $m_{\operatorname{\mathrm{p}red}(v)}(1-x_{v})$ multiplied by some monomial $m$. The only variable that can disappear in $M_{i+1}$ is $x_{v}$ and in this case the monomial $M_{i}=m_{\operatorname{\mathrm{p}red}(v)}x_{v}$. Therefore all the vertices in $\operatorname{\mathrm{p}red}(v)$ have pebbles on them and the pebble in $x_{v}$ can be removed. At the end of the refutation, when the 1 monomial is reached there are no pebbles left on $G.$ The number of pebbles present at any moment is the number of variables in any of the monomials and this is the degree of $\pi$. The number of pebbling steps needed is at most $d$ steps to place a pebble in each variable of $M_{1}=mx_{\rm sink}$ and then $t$ more pebbling steps. ∎
Observation 15.
@@
For the case of Polynomial Calculus it is known that strong degree lower bounds imply size lower bounds. If a set of unsatisfiable polynomials $P_{F}$ with $n$ variables and constant degree requires PC refutations of degree $s$, then any PC refutation of $P_{F}$ requires size at least $2^{\Omega(\frac{d^{2}}{n})}$ [IPS99]. The previous results show that this does not hold for Monomial Calculus. This follows from the fact that there are graph families $\{G_{n}\}_{n=0}^{\infty}$ with $n$ vertices and constant in-degree that require black pebbling space $\Omega(\frac{n}{\log n})$ [PTC77]. Theorem 14 implies that the pebbling formulas for this graph family needs degree $\Omega(\frac{n}{\log n})$. On the other hand, for every single-sink DAG with $n$ vertices there is a trivial black pebbling strategy using space $n$ and pebbling time $2n$. By Theorem 9 this implies that the pebbling formulas corresponding to the graphs in $\{G_{n}\}_{n=0}^{\infty}$ have MC refutations of quadratic size in $n$. This is a family of formulas with MC refutation degree $\Omega(\frac{n}{\log n})$ but having quadratic size refutations, a very different situation from the one in the PC case.
3.1 Degree separations
@@
The given relationships between MC and the black pebbling game allow for the immediate translation of pebbling results to Monomial Calculus. We start with some degree separations between MC and PC. The original motivation for introducing MC was the close connection between the degree complexity of the refutation of the graph isomorphism formulas in this proof system, and the Weisfeiler-Leman hierarchy [BG15]. Formulas corresponding to non-isomorphic graphs pairs that can only be distinguished using a large level of the WL algorithm, require a MC refutation with large degree. It was proven later in [AF23, GGPP19], that the degree of a PC refutation of the isomorphism formulas cannot be much smaller than in the MC case, in fact the degrees of a MC and a PC refutation can only be a constant factor apart. We improve this separation and obtain an almost optimal degree separations by considering the pebbling formulas. In [BCIP02] it was shown that pebbling formulas have constant PC degree and that for any directed acyclic graph $G$ with black pebbling price $B(G),$ the formula $\mathrm{Peb}(G)$ requires NS refutations with degree $\Omega(B(G))$. Since it is known that there are graph families $\{G_{n}\}_{n=0}^{\infty}$ with $\Theta(n)$ vertices and $B(G_{n})=\Omega(\frac{n}{\log n})$ [PTC77], this implies a degree separation of $\Omega(\frac{n}{\log n})$ between PC and NS. From Theorem 14 follows that this is in fact a degree separation between MC and PC.
@@
For the case of NS, from Theorem 9 and the equivalence between reversible pebbling price and NS degree from [dRMN+20], [dRMNR21], follows that a separation between reversible and black pebbling price for a graph family implies a degree separation between NS and MC for the corresponding pebbling formulas. For example it is known that a directed path graphs with $n$ vertices can be black pebbled with 2 pebbles but requires reversible pebbling number $\lceil\log n\rceil$ [Ben89]. Translated to pebbling formulas this means:
@@
There are other graph families for which a separation between the black and reversible pebbling prices by a logarithmic factor in the number of vertices is known, [CLNV15],[TW21]. The separations in pebbling for these graphs is translated into the next result.
@@
The question of whether the separation between reversible and black pebbling space can be larger than a logarithmic factor in the number of nodes is open. The best known degree separation between NS and and MC is slightly better. This was obtained in [GP17] with very different methods. Using a classic result from descriptive complexity [Imm81], the authors show that for for every constant $c\geq 1$ there are families of formulas $F_{n}$ with $O(n)$ variables that have a degree 3 MC refutation but require NS degree at least $\log^{c}({n}).$ It is also open whether this degree separation between NS and MC is optimal. Theorem 17 is the best known degree separation between NS and an MC refutation of degree 2.
3.2 Size-degree tradeoffs for MC
The close connections between black pebbling space and monomial calculus expressed in Theorems 9 and 14 make it possible to translate space-time tradeoffs for pebbling into degree-size tradeoffs for MC. There is a slight loss of the time parameter that comes from the extra space factor in the the MC refutation from Theorem 9. We present several such results as examples. The first one is an extreme tradeoff result that shows how decreasing the MC degree by one in a refutation can make the size increase exponentially.

### v2

@@ 3. Monomial Calculus and pebbling formulas
Proof 3.1.
It is convenient to consider here the equivalent notion of white pebbling. Let $\mathcal{P}=(\mathbb{P}_{0},\dots,\mathbb{P}_{t})$ be a white pebbling strategy for $G$ with $\mathbb{P}_{1}=\{z\}$ and $\mathbb{P}_{t}=\emptyset$ using $s$ pebbles. We show that for each pebbling configuration $\mathbb{P}_{i}$, $i\in[t]$, $\mathbb{P}_{i}=\{v_{i_{1}},\dots,v_{i_{k_{i}}}\}$ the monomial $m_{i}=\prod_{v\in\mathbb{P}_{i}}x_{v}$ can be derived from $\mathrm{Peb}_{G}$ and $m_{i-1}$ in degree $s$ and size 1 if $\mathbb{P}_{i}$ adds a pebble, or size 2s12s-1 if $\mathbb{P}_{i}$ removes a pebble. This proves the result since in the $t$ steps of the pebbling strategy half of the steps add a pebble and the other half of the steps remove a pebble (each added pebble has to be removed). The total number of steps is therefore $\frac{t}{2}+\frac{t}{2}(2s-1)=ts.$
@@
Pebble removal: If the configuration at pebbling step $i+1$ is reached after removing a white pebble from vertex $v$ and $\mathbb{P}_{i}=\{v,u_{i_{1}},\dots,u_{i_{k_{i}}}\}$ with $k_{i}\leq s-1$ then all predecessors of $v$ are in the set $\{u_{i_{1}},\dots,u_{i_{k_{i}}}\}.$ For the derivation of $m_{i+1}$ we can multiply the axiom $(1-x_{v})\prod_{u\in\operatorname{\mathrm{p}red}(v)}x_{u}$ by the variables in $\mathrm{Var}(m_{i})\setminus(\bigcup_{u\in\operatorname{\mathrm{p}red}(v)}x_{u}\cup\{x_{v}\})$, and add this polynomial to $m_{i}$ obtaining $m_{i+1}$. Since there are at most s1s-1 variables in $\mathrm{Var}(m_{i})\setminus(\bigcup_{u\in\operatorname{\mathrm{p}red}(v)}x_{u}\cup\{x_{v}\})$, the number of intermediate monomials added to the proof (counting also monomial $m_{i+1}$) is at most 2(s1)+1=2s1.2(s-1)+1=2s-1.
Observe that in all the steps in the refutation, at most two different monomials appear and the number of different variables in these monomials coincides with the largest of their degrees. This shows that the variable space of the MC refutation is bounded by $|\mathrm{Var}(m_{1})\cup\mathrm{Var}(m_{2})|\leq\max\{\deg(m_{1}),\deg(m_{2})\}$.
{obs}
The size bound $ts$ in the above proof comes from the way the MC rules are defined. As is the case of PC, in the multiplication rule only one variable at at time is multiplied, even when multiplying the axiom polynomials. When an axiom is multiplied by a monomial with several variables, all the intermediate polynomials contribute to the size of the MC refutation. This is different from the usual way to measure the size in the NS case, where intermediate monomials are not counted. If we would define the MC rules as those in NS, that is, if a whole monomial could be multiplied by an axiom in one step, the size of the MC proof would avoid the $s$ factor in the monomial size and obtain size $\frac{3t}{2}$ instead.
In order to prove a result in the other direction we consider a very restricted kind of refutation in MC, analogous to what is known as an input refutation in Resolution. In this kind of refutation in every Resolution step one of the parent clauses must be an axiom. Input Resolution is not complete, but it is complete for Horn formulas. We will show that the same is true for MC input refutations.
{defi}
A MC refutation $\pi$ of a contradictory set of polynomials $F$ is called an input refutation if there is a sequence of monomials $M_{0},\dots,M_{t}$ such that $M_{0}$ is the product of a monomial and an axiom, $M_{t}=1$ and for each $i$, $M_{i}$ is obtained by multiplying $M_{i-1}$ with a variable, or by the linear combination rule from $M_{i-1}$ and a monomial multiplied by an axiom polynomial. We will call the sequence of monomials $M_{0},\dots,M_{t}$ the backbone of the proof.
@@
Proof 3.2.
Let $d$ and $t$ be the degree and size of $\pi$. We can suppose that $\pi$ is multilinear. We prove the result by induction on $k$, the number of times the multiplication rule is applied to a monomial derived in $\pi$. In the base case $k=0$, $\pi$ is just a NS refutation of $P_{F}.$ This means that there is a linear combination of a set of polynomials $S$ that adds up to 1. Each of these polynomials has the form of a polynomial axiom multiplied by a monomial and since $F$ is a Horn formula, each polynomial in $S$ has either one or two monomials with coefficients in $\{1,-1\}$. We will represent such a polynomial $p=\alpha_{m}m+\alpha_{m^{\prime}}m^{\prime}$ by the pair of monomials $(m,m^{\prime})$. Since we are representing monomials by their set of variables, the monomial $1$ is represented by $\emptyset$. Clauses without positive literals are encoded as single monomials $(m)$. Some polynomial in $S$ has a single monomial otherwise the whole set $S$ would have a common root by setting all variables to 1. Moreover, we claim there has to be a sequence of polynomials $p_{0},\dots,p_{\ell}$ in $S$ represented by the monomials $(\emptyset,m_{1}),(m_{1},m_{2}),(m_{2},m_{3})\dots,(m_{\ell-1},m),(m).$ In order to see this, we define an undirected graph whose vertices are the different monomials in $S$ and there is an edge between two monomials $m,m^{\prime}$ if both appear in the same polynomial in $S$. In a minimal $\mathrm{MC}$ proof we only need to consider the connected component containing monomial $1$. Such component must have a vertex $v$ coming from a polynomial with a single monomial $m$ in $S$ because any set of binomials in $S$ can be satisfied by assigning all variables to 1. Therefore there is a path in the graph connecting vertices $1$ and $v$. Such a path with minimal length defines a sequence of edges $(1,v_{1}),(v_{1},v_{2}),\dots,(v_{\ell-1},v)$, and from it we can build the chain of polynomials $(\emptyset,m_{1}),(m_{1},m_{2}),(m_{2},m_{3})\dots,(m_{\ell-1},m),(m)$. This proves the claim.
@@
For the case $k>0$ let $m^{\prime}$ be the first monomial in the proof that is the result of a multiplication from a derived monomial $m$ and a variable $x$, $m^{\prime}=xm$ in $\pi$. The same argument as above shows that there is a sequence of polynomials $p_{0},\dots,p_{\ell},\hat{m}$ in $\pi$ from which we can extract an input monomial refutation that starts at $M_{0}=\hat{m}$ and derives at some point $M_{i}=m$. In the next step the multiplication rule is applied to obtain $M_{i+1}=m^{\prime}.$ Observe that the set of polynomials $m^{\prime}\cup P_{F}$ still has the Horn property and that there is a sub-proof of $\pi$ that derives the monomial 1 from this set applying the multiplication rule at most k1k-1 times. By induction hypothesis we know that there is a sequence of polynomials $p^{\prime}_{0},\dots,p^{\prime}_{\ell^{\prime}}$ in $\pi$ represented by the monomials $(\emptyset,m^{\prime}_{1}),(m^{\prime}_{1},m^{\prime}_{2}),\dots,(m^{\prime}_{r},m^{\prime})$ from which an input refutation of $P_{F}\cup m^{\prime}$ can be extracted. We can put together both input MC refutations $M_{0}\dots M_{i}$ and $M_{i+1},\dots,1$. Again we can assume that all the monomials in the backbone are different since otherwise we could shorten $\pi^{\prime}$. By the same argument as in the base case the size and degree of the input MC refutation cannot be larger than that of $\pi.$
@@
Because of Corollary 3 we can suppose that there exists an input MC refutation with monomials $M_{0},\dots M_{t}$ starting with $M_{1}=mx_{\rm sink}$ for some monomial $m$ and with $M_{t}=1$. We describe a strategy for a white pebbling of $G$ following $\pi$. As explained in the preliminaries, this is equivalent to a black pebbling strategy. At each step $i$ only the vertices corresponding to variables in $M_{i}$ have a pebble on them. In a multiplication step a new pebble is added, which is always possible in a white pebbling strategy. We only have to show that in case variables disappear when going from $M_{i}$ to $M_{i+1}$, this is also a correct pebbling move. But in this case, the step from $i$ to $i+1$ is a linear combination of $M_{i}$ with the axiom for some vertex $v$, $m_{\operatorname{\mathrm{p}red}(v)}(1-x_{v})$ multiplied by some monomial $m$. The only variable that can disappear in $M_{i+1}$ is $x_{v}$ and in this case the monomial $M_{i}=m_{\operatorname{\mathrm{p}red}(v)}x_{v}$. Therefore all the vertices in $\operatorname{\mathrm{p}red}(v)$ have pebbles on them and the pebble in $x_{v}$ can be removed. At the end of the refutation, when the 1 monomial is reached there are no pebbles left on $G.$ The number of pebbles present at any moment is the number of variables in any of the monomials and this is the degree of $\pi$. The number of pebbling steps needed is at most $d$ steps to place a pebble in each variable of $M_{1}=mx_{\rm sink}$ and then $t$ more pebbling steps.
{obs}
@@
For the case of Polynomial Calculus it is known that strong degree lower bounds imply size lower bounds. If a set of unsatisfiable polynomials $P_{F}$ with $n$ variables and constant degree requires PC refutations of degree $d$, then any PC refutation of $P_{F}$ requires size at least $2^{\Omega(\frac{d^{2}}{n})}$ [20]. The previous results show that this does not hold for Monomial Calculus. This follows from the fact that there are graph families $\{G_{n}\}_{n=0}^{\infty}$ with $n$ vertices and constant in-degree that require black pebbling space $\Omega(\frac{n}{\log n})$ [28]. Theorem 4 implies that the pebbling formulas for this graph family needs degree $\Omega(\frac{n}{\log n})$. On the other hand, for every single-sink DAG with $n$ vertices there is a trivial black pebbling strategy using space $n$ and pebbling time $2n$. By Theorem 1 this implies that the pebbling formulas corresponding to the graphs in $\{G_{n}\}_{n=0}^{\infty}$ have MC refutations of quadratic size in $n$. This is a family of formulas with MC refutation degree $\Omega(\frac{n}{\log n})$ but having quadratic size refutations, a very different situation from the one in the PC case.
3.1. Degree separations
@@
The given relationships between MC and the black pebbling game allow for the immediate translation of pebbling results to Monomial Calculus. We start with some degree separations between MC and PC. The original motivation for introducing MC was the close connection between the degree complexity of the refutation of the graph isomorphism formulas in this proof system, and the Weisfeiler-Leman hierarchy [9]. Formulas corresponding to non-isomorphic graphs pairs that can only be distinguished using a large level of the WL algorithm, require a MC refutation with large degree. It was proven later in [1, 16], that the degree of a PC refutation of the isomorphism formulas cannot be much smaller than in the MC case, in fact the degrees of a MC and a PC refutation can only be a constant factor apart. We improve this separation and obtain an almost optimal degree separations by considering the pebbling formulas. In [10] it was shown that pebbling formulas have constant PC degree and that for any directed acyclic graph $G$ with black pebbling price $B(G),$ the formula $\mathrm{Peb}(G)$ requires NS refutations with degree $\Omega(B(G))$. Since it is known that there are graph families $\{G_{n}\}_{n=0}^{\infty}$ with $\Theta(n)$ vertices and $B(G_{n})=\Omega(\frac{n}{\log n})$ [28], this implies a degree separation of $\Omega(\frac{n}{\log n})$ between PC and NS. From Theorem 4 follows that this is in fact a degree separation between MC and PC.
@@
For the case of NS, from Theorem 1 and the equivalence between reversible pebbling price and NS degree from [14], [15], follows that a separation between reversible and black pebbling price for a graph family implies a degree separation between NS and MC for the corresponding pebbling formulas. For example it is known that a directed path graphs with $n$ vertices can be black pebbled with 2 pebbles but requires reversible pebbling number $\lceil\log n\rceil$ [8]. Translated to pebbling formulas this means:
@@
There are other graph families for which a separation between the black and reversible pebbling prices by a logarithmic factor in the number of vertices is known, [11],[33]. The separations in pebbling for these graphs is translated into the next result.
@@
The question of whether the separation between reversible and black pebbling space can be larger than a logarithmic factor in the number of nodes is open. The best known degree separation between NS and and MC is slightly better. This was obtained in [17] with very different methods. Using a classic result from descriptive complexity [19], the authors show that for every constant $c\geq 1$ there are families of formulas $F_{n}$ with $O(n)$ variables that have a degree 3 MC refutation but require NS degree at least $\log^{c}({n}).$ It is also open whether this degree separation between NS and MC is optimal. Theorem 6 is the best known degree separation between NS and an MC refutation of degree 2.
3.2. Size-degree tradeoffs for MC
The close connections between black pebbling space and monomial calculus expressed in Theorems 1 and 4 make it possible to translate space-time tradeoffs for pebbling into degree-size tradeoffs for MC. There is a slight loss of the time parameter that comes from the extra space factor in the MC refutation from Theorem 1. We present several such results as examples. The first one is an extreme tradeoff result that shows how decreasing the MC degree by one in a refutation can make the size increase exponentially.